OpenStudy (anonymous):

What would the table of values look like for: 3x + 4y = 18 X + 3y = 11 With at least three rows :-)

OpenStudy (bahrom7893):

This is sooo simple... For the first one: x = -2, -1, 0, 1, 2 y = (plug those values in) Repeat the same thing for the 2nd one

OpenStudy (anonymous):

Is that a table of values for each linear equation?

OpenStudy (bahrom7893):

well for each separate..

OpenStudy (bahrom7893):

but you do the same thing

OpenStudy (anonymous):

Can you elaborate on what I have to do to get it for y?

OpenStudy (bahrom7893):

well to make your life easier get both equations in form y =

OpenStudy (bahrom7893):

3x + 4y = 18 4y = 18-3x y = (18-3x)/4

OpenStudy (bahrom7893):

All right I'll solve one out for you, next one you're doing on your own: X Y -2 (18-3*(-2))/4 -1 (18-3*(-1))/4 0 (18-3*(0))/4 1 (18-3*(1))/4 2 (18-3*(2))/4

OpenStudy (bahrom7893):

See, that's 5 rows... actually -1,0,1 would suffice

OpenStudy (anonymous):

I don't use fractions

OpenStudy (bahrom7893):

Well simplify them... im feeling too lazy for this, whip out your calculator and solve... first one: (18+6)/4 = 24/4 = 6

OpenStudy (bahrom7893):

still didn't finish? brizy come on this isn't an answer getting site... uhmm actually u can debate that lol